OpenStudy (aravindg):

find condition that one root of the equation ax^2+bx+c =0 may be n times the othr

OpenStudy (anonymous):

here , sum of root is \[\alpha +\beta= -b/a\]

OpenStudy (anonymous):

now , \[\alpha +n \alpha= -b/a\]

OpenStudy (anonymous):

\[\alpha(1+n)=-b/a\]

OpenStudy (anonymous):

so, \[\alpha= -b/a(1+n)\]

OpenStudy (anonymous):

now, \[\alpha \beta=c/a\]

OpenStudy (aravindg):

..........

OpenStudy (anonymous):

\[\alpha ^{2}=c/na\]

OpenStudy (anonymous):

so, condition is n= ac/(b^2-ac)